# Electronic structure and conductivity in functionalized multi-layer Black Phosphorene

### v1

@@ Abstract
Phosphorene and its components are highly reactive to oxygen when exposed to ambient conditions due to the presence of lone pairs of electrons on phosphorus atoms. Functionalization serves as a solution to prevent the chemical degradation of these materials. In this study, we investigate the impact of a rather strong covalent or non-covalent functionalization of phosphorene (monolayer Black Phosphorus (BP)), few-layer BP, and bulk BP. We use an effective tight-binding Hamiltonian that corresponds to one orbital per site, for which covalent functionalization is simulated by atomic vacancies, and non-covalent functionalization is simulated by Anderson disorder. We show that these two types of functionalisation act differently on the electronic structure and quantum diffusion, in particular on gap and mobility gap for a high degree of functionalisation.
@@ I Introduction
[…]
@@
Several recent experiments have been carried out recently to protect few-layer BP from chemical degradation due to environment, and to tailor its electrical properties. Covalent and non-covalent functionalization have been realized experimentally to enhance its environmental stability [12, 13, 14, 15, 16, 17]. The lone pair electrons of phosphorus atoms create a chemical bond with the adatoms. Additionally, theoretical investigations have contributed to develop the chemistry of few-layer BP using DFT models [18, 19, 20]. Yet crucial aspects of its reactivity remain largely unexplored and there is a lack of consensus between different interpretations of the experimental results. Functionalization can be realized on the surface of few-layer BP or of phosphorene but it can also be randomly distributed among the layers [21, 22]. It has also been demonstrated that functionalization can be covalent with resonant scatterers, such as functional groups containing free radicals like hydrogen, halogens, hydroxyl, thiol groups, and organic substituents which are important for biological applications [23, 11]. Functionalization can also be non covalent with non-resonant scatterers such as cesium, potassium or some organic molecules based on Van der Waals interaction [21, 24, 25, 26].
@@
In this paper, we focus our study on the electronic structure and the conductivity of functionalized monolayer, few-layer BP, and bulk by using the tight-binding (TB) model developed by Rudenko et al. [27, 28]. In the case of graphene, the description of its electronic properties at low energy is very simple because it is performed by the TB Hamiltonian involving only one hopping parameter. But for phosphorene, the use of this method is very complicated because it has four orbitals per atom and we must take into account the mixture of states of different symmetries. We model covalent functionalization using vacancies and non-covalent functionalization through Anderson disorder. We discuss the effects of these types of adsorbates on the density of States (DOS), conductivity, and mobility in phosphorene, few layers of black phosphorus (BP), and the bulk. Our idea is to compare the effect depending on the number of phosphorus layers, the concentration of adatoms, and the nature of the functionalization and try to emphasize the anisotropy of the material. In contrast to other theoretical research conducted through DFT calculations [3, 29, 30] and tight-binding Hamiltonian method [31], our research stands out by its unique emphasis on exploring transport properties. Specifically, we delve into the impacts of both covalent and non-covalent functionalization, particularly under conditions of high disorder where the Boltzmann semi-classical approach is not applicable.
@@
The paper is organized as follows. The atomic structure and the TB Hamiltonian are presented in Sec. II. In Sect. III, the numerical method for transport calculation is briefly detailed as well as the calculated transport coefficients. Results for resonant scatterers and non-resonant scatterers are presented in Sec. IV.1 and IV.2, respectively. Sec. V is dedicated to the discussion and conclusion. Additional figures are shown in the Supplemental Material [32].
@@ II Phosphorene structure and effective Hamiltonian
Figure 1: Atomic structure of bilayer Black Phosphorene (BP). The two plans have different colors. (left) Perspective side view, (right) Top view.
@@
In phosphorene, phosphorus atoms are joined by covalent bonds with bond angles of ${}^{\circ}$ and ${}^{\circ}$ [21, 22] (FIG. 1). The intralayer bonding results from the $sp^{3}$ hybridization of P atoms, giving rise to three bonding orbitals per two atoms augmented by lone pairs associated with each atom. It results in a strong stability in crystal structure because these angles are close to perfect tetragonal ones. The bonds out of the plane have an angle of approximately ${}^{\circ}$ with respect to the plane and are responsible for the puckered structure of phosphorene in contrast with graphene and 2D metal chalcogenides. The two first neighbor atoms in the plane are at the distances 2.224 Å, and the third neighbor one out of the plane is at 2.244 Å. The crystal structure of black phosphorus is orthorhombic with the lattice parameters 4.376, 3.314 and 10.478 Å for $a$, $b$, and $c$ respectively [33]. In the phosphorene, the primitive cell contains four atoms at positions $\pm$ $\left(ua,0,vc\right)$ and $\pm$ $\left((1/2-u)a,b/2,vc\right)$ with $u=0.08056$ and $v=0.10168$. The puckered honeycomb lattice formed by covalently bounded phosphorus atoms results in a significant in-plane anisotropy, similar to bulk phosphorus (BP), which distinguishes it from many other isotropic 2D materials [34, 35]. This anisotropy has implications for electronic properties of phosphorene such as an energy band structure, a density of states, a conductivity, etc… [36, 10, 37, 38]. The $x$-axis is the armchair direction and the $y$ one is the zigzag direction (FIG. 1).
@@
To study the effect of random disorder one needs to perform numerical calculations in a supercell containing several hundred thousand atoms. That can be only done in a framework of a tight binding (TB) Hamiltonian. However, in contrast to graphene which involves only one hopping parameter between $p_{z}$ orbitals, the phosphorene cell has four atoms with four orbitals per atom, $s$, $p_{x}$, $p_{y}$ and $p_{z}$, that need to be considered. The exact modeling of the TB Hamiltonian is a challenge because of the mixture of the orbital $s$ and orbitals $p$. However, it was been shown [39, 40, 27] that the $p_{z}$ orbital has a large contribution to the valence and conduction band around the band gap and its role is predominant compared to the roles of the other orbitals $s$, $p_{x}$ and $p_{y}$. Therefore, following Rudenko et al. [27, 28], it is possible to consider only one orbital per atom that looks like a $p_{z}$ without being exactly the $p_{z}$ orbital. This approximation is reasonable as long as the energy is fairly close to the gap.
@@
Indeed this effective model has been successfully used to describe the studies that concern phosphorene nanoribbons [41, 42, 43], electric [41, 44] and magnetic fields [45, 31, 46], field-effect electronic devices [5] and quantum dots [47]. We also use that effective TB model which corresponds to one orbital $|i\rangle$ per P site [28],
@@
where $i$ is the orbital located at $\bf{r}_{i}$ and the sum runs over neighboring $i$, $j$ lattices sites. $\epsilon_{i}$ is the on-site energy of the electron and it takes the same value for all sites because we suppose that all sites are equivalent. $t^{\parallel}_{ij}$ ($t^{\perp}_{ij}$) is the intralayer (interlayer) hopping parameter between the $i$ and $j$ sites. Those parameters are obtained [28] by the construction of a set of four maximally localized Wannier functions corresponding to $p_{z}$-like orbitals from DFT calculation with quasiparticle GW approximation [28]. Typically, we use the TB parameters proposed by Rudenko et al. in Ref. [28]: ten intralayer and four interlayer hopping parameters over the distances between the corresponding interacting sites up to 5.5 Å.
FIG. 2 shows the total densities of states (DOS) for multilayer of BP (from monolayer to bulk). One can see that the band gap of the monolayer is about 1.75 eV. The origin of this gap is the fact that this TB Hamiltonian involved the nearest neighbor atoms, two in-plane hopping terms $t_{1}=-1.22$ eV and one out-of-plane hopping term $t_{2}=3.67$ eV. The distinct sign of $t_{2}$ in comparison to $t_{1}$ implies the opening of the gap [27, 28, 48]. It should be noted that more theoretical works have predicted a value energy gap ranging from $1$ eV [6, 10, 49] up to $2$ eV [9] depending on the calculation method, and the measured value experimentally is around 2 eV [50, 51].
The band gap decreases according to the number of layers and for the bulk the calculated value is around 0.3 eV. A step-like behavior is the signature of reducing dimensionally.
Figure 2: Tight binding (TB) density of states (DOS) of the phosphorene, few-layer BP (up to 5 layers), and bulk, calculated from TB parameters proposed by Ref. [28].
@@ III Numerical method for transport
The electronic properties of the multilayer of phosphorene are calculated by using the real space Kubo Greenwood formalism computed by recursion and polynomial expansion method [52, 53, 54, 55, 56]. It allows us to compute efficiently quantum transport in a large system for which calculation by diagonalization is not numerically feasible. This formalism has been used to study the effect of vacancies, defects and disorder on quantum transport in different materials such as graphene monolayer and bilayer [57, 58, 59, 60, 61, 62, 63], carbon nanotube [64, 65, 66], quasicristals [56], perovskites [67], etc…
This real space method allows the include local defects directly into the Hamiltonian, and it allows a random distribution of defects in the huge supercell. To make a systematic study of the effect of defects on the electronic properties, we study two types of generic defects:
(i) Resonant defect, i.e. defect that creates a covalent bond with a P atom of the phosphorene. As the system is described by the effective four-band Hamiltonian, with one $p_{z}$-like orbital per phosphorus atom, we use the same approach as in the case of graphene to model resonant scatterers [59, 57, 58]. Indeed an adatom or ad-molecule bound covalently with the P atom is almost equivalent for the electron bands to remove the corresponding orbitals of the bounded P. Thus a resonant scatter is modelized by a vacant P atom. In our calculation defects (vacancies) are randomly over the whole supercell of 640 000 P atoms per layer.
(ii) Non-resonant defect, i.e. non-covalent functionalization, such as Van der Waals bonding, that creates a small local perturbation of the electronic structure. A common method of dealing with this type of defect is to use the Anderson disorder where the on-site P energies $\epsilon_{i}$ (equation (1)) are randomly distributed in $[\epsilon_{i}^{0}-W/2,\epsilon_{i}^{0}+W/2]$. $W$ is the magnitude of that disorder and $\epsilon_{i}^{0}$ the on-site energy of the orbital $i$ without disorder from Ref. [28].
To evaluate the effect of the anisotropy of phosphorene on the conductivity $\sigma$, we compute [52, 53, 54, 55, 56] $\Delta X^{2}(E,t)$ and $\Delta Y^{2}(E,t)$, which are the average square of the quantum spreading along the armchair direction and the zigzag direction, respectively (FIG. 1), after a time $t$ and for states at energy $E$.
In such a calculation, all quantum effects –including all multiple-scattering effects– are taken into account to calculate the average square spreading $\Delta X^{2}$ without inelastic scattering i.e. at zero temperature. At finite temperature $T$, the inelastic scattering caused by the electron-phonon interactions is implanted by using the Relaxation Time Approximation (RTA)[59]. The conductivity in the $x$-direction ($y$-direction) is thus given by,
@@
where $E_{F}$ is the Fermi energy, $\tau_{i}$ is the inelastic scattering time, $n(E)$ is the total density of states (total DOS), $D$ the diffusivity along the $x$-axis ($y$-axis), and $L_{i}$ is the inelastic mean free path. $L_{i}(E_{F},\tau_{i})$ is the typical distance of propagation during the time interval $\tau_{i}$ for electrons at energy $E$. Roughly speaking $\tau_{i}$ is the time beyond which the velocity autocorrelation function goes exponentially to zero [59].
@@
Figure 3: With resonant scatterers (vacancies): (a) DOS for $c=1\,\%$ depending with number of layers for phosphorene, (b) $E_{g}$ for different concentrations and for 1, 2, 3, 4, 5 layers of phosphorene and bulk, (c) DOS versus concentration for monolayer, (d) DOS versus concentration for trilayer. $G_{0}=2e^{2}/h$.
@@
Figure 4: Conductivity at $T=300$ K (Eq. 6) per layer with resonant scatterers (vacancies): (a) $\sigma_{xx}$ versus $N_{e}$ for 1, 2, 3, 4, 5 layers and bulk and for $c=1$, (b) $\sigma_{xx}$ for monolayer (bold line) and trilayer (dashed line) for different concentrations, (c) $\sigma_{xx}/\sigma_{yy}$ versus $N_{e}$ for $c=1$, (d) $\sigma_{xx}/\sigma_{yy}$ versus $N_{e}$ for monolayer (bold line) and trilayer (dashed line), (e) Mobility versus $N_{e}$ for 1, 2, 3, 4, 5 layers and bulk and for $c=1\%$, (f) Mobility for (bold line) monolayer and (dashed line) trilayer for different concentrations $c$. $N_{e}$ is the density of charge carriers per layer. $G_{0}=2e^{2}/h$.
@@
Figure 5: With non-resonant scatterers (Anderson disorder): (a) Density of states for $W=1$ depending on the number of layers for phosphorene, (b) $E_{g}$ for different Anderson disorders and for 1, 2, 3, 4 layers of phosphorene and bulk. Dos for different $W$ values in (c) monolayer and (d) trilayer. $G_{0}=2e^{2}/h$.
@@
Figure 6: Conductivity at $T=300$ K (Eq. 6) per layer with non-resonant scatterers (Anderson disorder): (a) $\sigma_{xx}$ versus $N_{e}$ for 1, 2, 3, 4, 5 layers and bulk and for Anderson disorder $W=1$, (b) $\sigma_{xx}$ for monolayer and bilayer for different magnitude $W$ of disorders, (c) $\sigma_{xx}/\sigma_{yy}$ versus $N_{e}$ for $W=1$, (d) $\sigma_{xx}/\sigma_{yy}$ versus $N_{e}$ for monolayer and trilayer, (e) Mobility versus $N_{e}$ for 1, 2, 3, 4, 5 layers and bulk and for $W=2$, (f) Mobility for (bold line) monolayer and (dashed line) trilayer for different Anderson disorder. $N_{e}$ is the density of charge carriers per layer. $G_{0}=2e^{2}/h$.
@@
As the monolayer and its components are semiconductors, we use the Gaussian broadening of the spectrum to compute the DOS avoiding the tail expansion of the Lorentzian broadening [68]. In the present work, a Gaussian broadening is used with the Gaussian standard deviation of 10 meV. The microscopic conductivity $\sigma^{\rm m}(E)$ at energy $E$ is the maximum value of the conductivity,
$\sigma^{\rm m}(E)={\rm Max}\{\sigma(E)\}_{\tau_{i}},$ (5)
@@
calculated from ${\rm Max}\{D(E,\tau_{i})\}_{\tau_{i}}$. The microscopic conductivity at energy $E$ (Eq. 5) therefore includes the effects of temperature on the diffusivity of charge carriers for high temperature or room temperature [59], but it does not include the mixing of states of various energy $E$ due to the Fermi Dirac distribution. Thus, we calculate the conductivity at room temperature $\sigma$ as follows,
$\sigma(\mu_{c},T\simeq 300{\rm K})=-\int_{-\infty}^{+\infty}\sigma^{\rm m}(E)\frac{\partial f(\mu_{c})}{\partial E}{\rm d}E,$ (6)
@@
where $\mu_{c}$ is the chemical potential, $\sigma^{\rm m}(E)$ is calculated from (5), and $f$ is the Fermi-Dirac distribution function. At temperature $T$, the mobility $\mu$ is related with conductivity $\sigma$ by,
@@
where $N_{e}$ is the density of charge carriers with respect to the neutral case. Thus at room temperature ($T\simeq 300$ K):
$\mu(\mu_{c},T)=\frac{-1}{e\lvert N_{e}\rvert}\int_{-\infty}^{+\infty}\sigma_{M}(E)\frac{\partial f(\mu_{c})}{\partial E}{\rm d}E,$ (8)
@@
$N_{e}(\mu_{c},T)=\frac{1}{S}\int_{-\infty}^{+\infty}n(E)f(E,\mu_{c},T){\rm d}E-N_{e}^{0}$ (9)

### v2

# Electronic structure and conductivity in functionalized multilayer black phosphorene

@@ Abstract
Phosphorene and its components are highly reactive to oxygen when exposed to ambient conditions due to the presence of lone pairs of electrons on phosphorus atoms. Functionalization serves as a solution to prevent the chemical degradation of these materials. In this paper, we investigate the impact of relatively strong covalent or noncovalent functionalization on phosphorene (monolayer black phosphorus (BP)), few-layer BP, and bulk BP. We use an effective tight-binding Hamiltonian that corresponds to one orbital per site, wherein covalent functionalization is simulated by atomic vacancies, and noncovalent functionalization is simulated by Anderson disorder. We demonstrate that these two types of functionalization act differently on the electronic structure and quantum diffusion, particularly affecting the gap and mobility characteristics, especially with a high degree of functionalization. However, we also show that the mobility gap is not significantly modified by the two types of defect. We also analyze the electron-hole asymmetry that is more important for multilayer and bulk BP.
preprint: Phys. Rev. B 110, 045150 (2024)
@@ I Introduction
The scientific community has shown keen interest in two-dimensional (2D) materials, focusing on their potential applications in electronics, sensing, dye degradation, and various biomedical uses [1, 2, 3, 4]. An initial 2D crystal realized by mechanical exfoliation was graphene. It is an interesting material for its outstanding electronic and mechanical properties. However, its zero band gap limited its use in optoelectronics and semiconductor devices. Therefore, several research studies have been carried out to find other 2D materials with a band gap and large carrier mobility to replace graphene. In 2014, a successful exfoliation of few-layer black phosphorus (BP), the most stable allotrope of phosphorus, was realized, and phosphorene, i.e., monolayer black phosphorus, was obtained [5, 6, 7]. In contrast to graphene, phosphorene is a 2D semiconductor with a direct band gap with a high hole mobility up to 1000 cm2V-1s-1 at room temperature [8]. Its band gap, $E_{g}$, varies from 2 eV [9] for phosphorene monolayer to 0.3 eV for the bulk [10, 11]. Monolayer and multilayer BPs exhibit a unique in-plane structural anisotropy which reflects in the band anisotropy around the gap and hence in the effective masses of electrons and holes [10, 12, 7, 13].
Functionalization or disorder can modify the band gap value to tailor these materials for desired applications. Phosphorene and its components are highly sensitive to the environment. This sensitivity arises from their pronounced reactivity to oxygen under ambient conditions, which may constrain their practical use in real devices. Therefore, it is crucial to investigate the impact of defects and impurities on the electronic structure and transport properties for this purpose [14].
@@
Several recent experiments have been carried out to protect few-layer BP from chemical degradation due to the environment and tailor its electrical properties. Covalent and noncovalent functionalization have been realized experimentally to enhance its environmental stability [15, 16, 17, 18, 19, 20]. The lone pair electrons of phosphorus atoms create a chemical bond with the adatoms. Additionally, theoretical investigations have contributed to the development of the chemistry of few-layer BP using DFT models [21, 22, 23]. Yet, crucial aspects of its reactivity remain largely unexplored, and there is a lack of consensus between different interpretations of the experimental results. Functionalization can be carried out on the surface of few-layer BP or that of phosphorene, but can also be randomly distributed among the layers [24, 25]. It has also been demonstrated that functionalization can be covalent, with resonant scatterers, such as functional groups containing free radicals like hydrogen, halogens, hydroxyl, thiol groups, and organic substituents, which are important for biological applications [26, 14]. Functionalization can also be noncovalent, with nonresonant scatterers such as cesium, potassium, or some organic molecules based on van der Waals interaction [24, 27, 28, 29].
@@
In this paper, we focus our study on the electronic structure and the conductivity of functionalized monolayer, few-layer BP, and bulk by using the tight-binding (TB) model developed by Rudenko et al. [30, 31]. In the case of graphene, the description of its electronic properties at low energy is very simple because it is performed by the TB Hamiltonian involving only one hopping parameter. However, for phosphorene, using this method is very complicated because it has four orbitals per atom, and we must take into account the mixture of states of different symmetries. With the Rudenko et al. simplified Hamiltonian, we model covalent functionalization using vacancies and noncovalent functionalization through Anderson disorder. We discuss the impact of these types of adsorbates on the density of states (DOS), conductivity, and mobility in phosphorene, few-layer BP, and the bulk. Our idea is to compare the effect depending on the number of phosphorus layers, the concentration of adatoms, and the nature of the functionalization and try to emphasize the anisotropy of the material. In contrast to other theoretical research conducted through DFT calculations [32, 33, 3, 34] and the TB Hamiltonian method [35], our research stands out by its unique emphasis on exploring transport properties. Specifically, we delve into the impact of both covalent and noncovalent functionalization, particularly under conditions of high disorder where the Boltzmann semi-classical approach is not applicable.
@@
The paper is organized as follows: The atomic structure and the TB Hamiltonian are presented in Sec. II. In Sect. III, the numerical method for transport calculation is briefly detailed as well as the calculated transport coefficients. Results for resonant scatterers and nonresonant scatterers are presented in Sec. IV.1 and IV.2, respectively. Section V is dedicated to the discussion and conclusion. Additional figures are shown in the Supplemental Material [36] (see Appendix Sec. VI).
@@ II Phosphorene structure and effective Hamiltonian
Figure 1: Atomic structure of bilayer black phosphorene (BP). The two planes have different colors. Left: Perspective side view, Right: Top view.
@@
In phosphorene, phosphorus atoms are joined by covalent bonds with bond angles of 96.34 and 103.09 [24, 25] (Fig. 1). The intralayer bonding results from the $sp^{3}$ hybridization of P atoms, giving rise to three bonding orbitals per two atoms augmented by lone pairs associated with each atom. This results in strong stability in the crystal structure because these angles are close to perfect tetragonal ones. The bonds out of plane have an angle of approximately 45 degrees with respect to the plane and are responsible for the puckered structure of phosphorene, in contrast with graphene and 2D metal chalcogenides. The first two neighbor atoms in the plane are at distances of 2.224 Å, and the third neighbor out of plane is at 2.244 Å. The crystal structure of black phosphorus is orthorhombic with lattice parameters of 4.376 Å, 3.314 Å, and 10.478 Å for $a$, $b$, and $c$, respectively [37]. In phosphorene, the primitive cell contains four atoms at positions $\pm$ $\left(ua,0,vc\right)$ and $\pm$ $\left((1/2-u)a,b/2,vc\right)$ with $u=0.08056$ and $v=0.10168$. The puckered honeycomb lattice formed by covalently bonded phosphorus atoms results in significant in-plane anisotropy, similar to bulk BP, distinguishing it from many other isotropic 2D materials [12, 38]. This anisotropy has implications for electronic properties of phosphorene such as energy band structure, conductivity, etc. [39, 10, 40, 41]. The $x$ axis is the armchair direction, and the $y$ axis is the zigzag direction (Fig. 1).
@@
To study the effect of random disorder, one needs to perform numerical calculations in a supercell containing several hundred thousand atoms. That can only be done in the framework of a TB Hamiltonian. However, in contrast to graphene which involves only one hopping parameter between $p_{z}$ orbitals, the phosphorene cell has four atoms with four orbitals per atom, $s$, $p_{x}$, $p_{y}$, and $p_{z}$, that need to be considered. The exact modeling of the TB Hamiltonian is a challenge because of the mixture of the orbital $s$ and orbitals $p$. However, it has been shown [42, 43, 30] that the $p_{z}$ orbital has a large contribution to the valence and conduction band around the band gap, and its role is predominant compared to the roles of the other orbitals $s$, $p_{x}$, and $p_{y}$. Therefore, following Rudenko et al. [30, 31], it is possible to consider only one orbital per atom that looks like a $p_{z}$ without being exactly the $p_{z}$ orbital. This approximation is reasonable as long as the energy is fairly close to the band gap.
@@
Indeed, this effective model has been successfully used to describe studies concerning phosphorene nanoribbons [44, 45, 46], electric [44, 47] and magnetic fields [48, 35, 49], field-effect electronic devices [5], and quantum dots [50]. We also use the effective TB model which corresponds to one orbital $|i\rangle$ per P site [31],
@@
where $i$ is the orbital located at $\bf{r}_{i}$ and the sum runs over neighboring $i$, $j$ lattice sites. $\epsilon_{i}$ is the on-site energy of the electron and it takes the same value for all sites because we suppose that all sites are equivalent. $t^{\parallel}_{ij}$ ($t^{\perp}_{ij}$) is the intralayer (interlayer) hopping parameter between the $i$ and $j$ sites. Those parameters are obtained [31] by the construction of a set of four maximally localized Wannier functions corresponding to $p_{z}$-like orbitals from DFT calculation with quasiparticle $GW$ approximation [31]. We use the TB parameters proposed by Rudenko et al. in Ref. [31]: ten intralayer and four interlayer hopping parameters over distances between the corresponding interacting sites up to 5.5 Å.
Figure 2 shows the total DOS of defect-free multilayers of BP (from monolayer to bulk). One can see that the band gap of the monolayer is about 1.75 eV. The origin of this gap is the fact that this TB Hamiltonian involved the nearest neighbor atoms, two in-plane hopping terms $t_{1}=-1.22$ eV, and one out-of-plane hopping term $t_{2}=3.67$ eV. The distinct sign of $t_{2}$ in comparison to $t_{1}$ implies the opening of the gap [30, 31, 51]. It should be noted that more theoretical works have predicted an energy gap value ranging from 1 eV [6, 10, 52] up to 2 eV [9] depending on the calculation method, and the measured value experimentally is around 2 eV [11, 53, 54].
As already reported experimentally and theoretically [55, 56], the band gap decreases with the number of layers, and for the bulk, its value is around 0.3 eV. This band gap evolution is consistent with an increase of the transversal confinement when the number of layers decreases. It is also interesting to note that as the number of layers decreases, the electron-hole asymmetry becomes less pronounced. Indeed, for the monolayer, the DOS is almost symmetric with respect to the band gap. Whereas, as the number of layers increases, the DOS of the valency band becomes lower than the DOS of the conduction band (see inset of Fig. 2).
Figure 2: Tight binding (TB) density of states (DOS) of defect-free phosphorene, few-layer BP (up to 5 layers), and bulk BP, calculated using the TB parameters proposed by Ref. [31].
@@ III Numerical method for transport
The electronic properties of the multilayer phosphorene are calculated using the real-space Kubo-Greenwood formalism computed by recursion and polynomial expansion method [57, 58, 59, 60, 61]. This approach efficiently computes quantum transport in large systems for which diagonalization is not numerically feasible. The formalism has been employed to study the effect of vacancies, defects, and disorder on quantum transport in various materials, including monolayer and bilayer graphene [62, 63, 64, 65, 66, 67, 68], carbon nanotubes [69, 70, 71], quasicrystals [61], and perovskites [72].
This real-space method enables the inclusion of local defects directly into the Hamiltonian and allows for a random distribution of defects in the huge supercell. To systematically study the effect of defects on electronic properties, we investigate two types of generic defects:
(i) Resonant defects, i.e. defects that create a covalent bond with a P atom of phosphorene. As the system is described by the effective four-band Hamiltonian, with one $p_{z}$-like orbital per phosphorus atom, we employ the same approach as in the case of graphene to model resonant scatterers [64, 62, 63]. An adatom or ad-molecule bound covalently with the P atom is almost equivalent for the electron bands to removing the corresponding orbitals of the bonded P. Thus, a resonant scatter is modeled by a vacant P atom. In our calculations, defects (vacancies) are randomly distributed over the entire supercell of 640000 P atoms per layer.
(ii) Nonresonant defects, i.e. noncovalent functionalization, such as van der Waals bonding, that create a small local perturbation of the electronic structure. A common method for dealing with this type of defect is to use Anderson disorder, where the on-site energies $\epsilon_{i}$ of all sites (Eq. (1)) are randomly distributed in $[\epsilon_{i}^{0}-W/2,\epsilon_{i}^{0}+W/2]$. This uniform Anderson disorder is one of the simplest models (with only one parameter) that is commonly used for nonresonant defects [73]. Here, $W$ is the magnitude of that disorder, and $\epsilon_{i}^{0}$ is the on-site energy of the orbital $i$ without disorder from Ref. [31].
To evaluate the effect of the anisotropy of phosphorene on the conductivity $\sigma$, we compute [57, 58, 59, 60, 61] $\Delta X^{2}(E,t)$ and $\Delta Y^{2}(E,t)$, which represent the average square of the quantum spreading along the armchair direction and the zigzag direction, respectively (see Fig. 1), after a time $t$ and for states at energy $E$.
In such calculations, all quantum effects, including all multiple-scattering effects, are taken into account to calculate the average square spreading $\Delta X^{2}$ without inelastic scattering, i.e., at zero temperature. At finite temperature $T$, the inelastic scattering caused by the electron-phonon interactions is accounted for using the relaxation time approximation (RTA) [64]. The conductivity in the $x$-direction ($y$-direction) is given by,
@@
where $E_{F}$ is the Fermi energy, $\tau_{i}$ is the inelastic scattering time, $n(E)$ is the total DOS, $D$ is the diffusivity along the $x$ axis ($y$ axis), and $L_{i}$ is the inelastic mean free path. $L_{i}(E_{F},\tau_{i})$ represents the typical distance of propagation during the time interval $\tau_{i}$ for electrons at energy $E$. In simple terms, $\tau_{i}$ is the time beyond which the velocity autocorrelation function exponentially approaches zero [64].
@@
Figure 3: With resonant scatterers (vacancies): (a) DOS for defect concentration $c=1\,\%$ depending on the number of layers for phosphorene, (b) $E_{g}$ for different $c$ and for 1–5 layers of phosphorene and bulk, (c) monolayer DOS for different $c$, (d) trilayer DOS for different $c$. (e) (f) Microscopic conductivity (Eq. (5)) times defect concentration $c$, (bold line) $c\,\sigma^{\rm m}_{xx}$ and (dashed line) $c\,\sigma^{\rm m}_{yy}$, versus $E$, in (e) monolayer and (f) trilayer. These results show that $c\,\sigma^{\rm m}$ is almost independent of $c$ for energies outside the gap, which proves Boltzmann-type behavior. $G_{0}=2e^{2}/h$.
@@
Figure 4: Conductivity at $T=300$ K (Eq. (6)) per layer with resonant scatterers (vacancies): (a) $\sigma_{xx}$ versus $N_{e}$ for 1–5 layers and bulk and for $c=1$, (b) $\sigma_{xx}$ for the monolayer (bold line) and trilayer (dashed line) for different defect concentrations $c$, (c) $\sigma_{xx}/\sigma_{yy}$ versus $N_{e}$ for $c=1$, (d) $\sigma_{xx}/\sigma_{yy}$ versus $N_{e}$ for the monolayer (bold line) and trilayer (dashed line), (e) Mobility versus $N_{e}$ for 1–5 layers and bulk and for $c=1\%$, (f) Mobility for the monolayer (bold line) and trilayer (dashed line) for different concentrations $c$. $N_{e}$ is the density of charge carriers per layer. $G_{0}=2e^{2}/h$.
@@
Figure 5: With nonresonant scatterers (Anderson disorder): (a) Density of states for $W=1$ eV depending on the number of layers for phosphorene. (b) $E_{g}$ for different Anderson disorders and for 1–5 layers of phosphorene and bulk. DOS for different $W$ (eV) values in (c) the monolayer and (d) the trilayer. (e) (f) Microscopic conductivity (Eq. (5)) for different Anderson disorders: (bold line) $\sigma^{\rm m}_{xx}$ and (dashed line) $\sigma^{\rm m}_{yy}$ versus $E$, in (e) the monolayer and (f) trilayer. $G_{0}=2e^{2}/h$.
@@
Figure 6: Conductivity at $T=300$ K (Eq. (6)) per layer with nonresonant scatterers (Anderson disorder): (a) $\sigma_{xx}$ versus $N_{e}$ for 1–5 layers and bulk, and for Anderson disorder $W=1$ eV, (b) $\sigma_{xx}$ for monolayer and bilayer for different magnitudes $W$ (eV) of disorders, (c) $\sigma_{xx}/\sigma_{yy}$ versus $N_{e}$ for $W=1$ eV, (d) $\sigma_{xx}/\sigma_{yy}$ versus $N_{e}$ for monolayer and trilayer, (e) Mobility versus $N_{e}$ for 1–5 layers and bulk, and for $W=2$ eV, (f) Mobility for monolayer (bold line) and trilayer (dashed line) for different Anderson disorders. $N_{e}$ is the density of charge carriers per layer. $G_{0}=2e^{2}/h$.
@@
As the monolayer and its components are semiconductors, we use Gaussian broadening of the spectrum to compute the DOS, avoiding the tail expansion of Lorentzian broadening [74]. In the present paper, a Gaussian broadening with a standard deviation of $10$ meV is used.
The microscopic conductivity $\sigma^{\rm m}(E)$ at energy $E$ is defined as the maximum value of the conductivity, calculated from ${\rm Max}\{D(E,\tau_{i})\}_{\tau_{i}}$:
$\sigma^{\rm m}(E)={\rm Max}\{\sigma(E)\}_{\tau_{i}}.$ (5)
@@
The microscopic conductivity at energy $E$ (Eq. (5)) includes the effects of temperature on the diffusivity of charge carriers for high temperatures or room temperature [64], but it does not take into account the mixing of states of various energies $E$ due to the Fermi-Dirac distribution. Therefore, we calculate the conductivity at room temperature $\sigma$ as follows:
$\sigma(\mu_{c},T\approx 300\,{\rm K})=-\int_{-\infty}^{+\infty}\sigma^{\rm m}(E)\frac{\partial f(\mu_{c})}{\partial E}{\rm d}E,$ (6)
@@
where $\mu_{c}$ is the chemical potential, $\sigma^{\rm m}(E)$ is calculated from Eq. 5, and $f$ is the Fermi-Dirac distribution function.
At temperature $T$, the mobility $\mu$ is related to the conductivity $\sigma$ by:
@@
where $N_{e}$ is the density of charge carriers with respect to the neutral case at $T=0$. Thus, at room temperature ($T\approx 300$ K):
$\mu(\mu_{c},T)=\frac{-1}{e\lvert N_{e}\rvert}\int_{-\infty}^{+\infty}\sigma^{\rm m}(E)\frac{\partial f(\mu_{c})}{\partial E}{\rm d}E,$ (8)
@@
$N_{e}(\mu_{c},T)=\frac{1}{S}\int_{-\infty}^{+\infty}n(E)f(E,\mu_{c},T){\rm d}E-N_{e}^{0},$ (9)